# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v3

@@ 5 Datasets
This section provides an overview of the existing methods for generating and manipulating quantum data to be usable for QML tasks. After discussing the data pre-processing methods and the data encoding techniques, we present a list of quantum datasets.
5.1 Quantum Data
Some examples of quantum data can be:
  • Discrete data (binary or integer values).
  • Real continuous data (floating-point values).
  • Complex continuous data (complex values).
5.2 Motivation for QML Datasets
@@ 5.3 Data Pre-Processing
Since QML has not yet extensively demonstrated its advantages compared to classical ML, it is more susceptible to adoption by the industry. The current quantum computers have few qubits to test and are noisy, making it difficult to demonstrate the current and potential quantum advantage of QML methods. Utilizing quantum methods for ML in the pre-processing phase, it is possible to achieve better classical encoding and performance of quantum classifiers (Mancilla & Pere, 2022). Since many proposed QML applications rely on using well-known datasets, especially in the context of quantum data with constraints generated due to the underlying quantum mechanics for representing it as quantum states, it is important to formulate and standardize its collection and encoding processes over time to maintain quality (Sierra-Sosa et al., 2021).
@@ 5.5 Quantum Datasets Collection
  • The QDataSet (Perrier et al., 2022), released in 2022, is composed of $52$ datasets based on simulations of one-qubit and two-qubit systems evolving in the presence or absence of noise. It provides a large-scale set of datasets for QML practitioners to train, benchmark, and develop classical and quantum algorithms for common tasks in quantum sciences, such as quantum control, quantum tomography, and noise spectroscopy. It has been generated using customized code running on base-level Python packages to facilitate interoperability and portability across common QML platforms. Each dataset consists of $10,000$ samples and includes a range of information (stored in list, matrix, or tensor format) regarding quantum systems and their evolution, such as quantum state vectors, drift and control Hamiltonians, unitaries, Pauli measurement distributions, time series data, pulse sequence data for square and Gaussian pulses, noise and distortion data. The total compressed size of the QDataSet (compressed with Pickle and zip) is around $14$ TB, while its uncompressed size is around $100$ TB.
  • The QM7, QM8, and QM9 datasets are subsets of the GDB-13 (Blum & Reymond, 2009) and GDB-17 (Ruddigkeit et al., 2012) databases, which contain billions of organic molecules. The QM7 (Rupp et al., 2012) is composed of $7,165$ molecules of up to $23$ atoms (including $7$ heavy atoms such as C, N, O, and S), represented using the Coulomb matrix. An extension of the QM7 dataset for multitask learning, called QM7B (Montavon et al., 2013), has $13$ additional properties (e.g., polarizability, HOMO and LUMO eigenvalues, excitation energies) and contains a total of $7,211$ molecules. The QMB8 (Ramakrishnan et al., 2015) has a training set of $10,000$ molecules with up to $8$ heavy atoms. The QM9 (Ramakrishnan et al., 2014) contains $133,885$ molecules with up to $9$ heavy atoms. These databases serve as benchmarks for hybrid QML models and systematic identification of structure-property relationships.
@@
  • The PennyLane Datasets (Bergholm et al., 2018) are composed of the Quantum Chemistry (QChem) Datasets for common molecular systems and the Quantum Many-Body Physics (QMBP) Datasets. The QChem datasets contain the electronic structure data for $42$ different geometries of molecules such as linear hydrogen chains, metallic and non-metallic hydrides, and charged species. For each geometry of the molecule, it is possible to extract molecular data (i.e., information regarding the molecule), Hamiltonian data (i.e., Hamiltonian for the molecular system), tapering data (i.e., features based on Z2 symmetries of the molecular Hamiltonian for performing tapering), tapered observables data, and variational data. The QMBP datasets contain spin models displaying quantum correlations. For each spin system, datasets are generated for $1$-D lattices (linear chain) and $2$-D lattices (rectangular grid). Each dataset contains values for $100$ different configurations of tunable parameters such as the external magnetic field and coupling constants.
  • The TensorFlow Quantum Data builds upon classical datasets such as MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017). It implements the method proposed by Farhi & Neven (2018) to use binary encoding for representing each pixel with a qubit, with the state depending on the pixel’s value. According to the experimental results of Huang et al. (2021), this method achieves quantum speedup in the fault-tolerant regime compared to the classical system.
@@
  • The Mendley Datasets for Quantum Circuits Mapping (Acampora et al., 2021) helps to solve the circuit mapping problem as a classification task. Each dataset contains features related to the calibration data of the physical device and others related to the generated quantum circuit for specific IBM quantum machines, such as IBMQ Santiago, IBMQ Athens, and IBMQ 16 Melbourne.
@@
  • The QFlow Lite dataset (Zwolak et al., 2018) contains a Python-based software suite to train neural networks to recognize the state of a device and differentiate between states in experimental data. It consists of $1,001$ idealized simulated measurements with gate configurations sampling (stored as $100\times 100$ pixel maps) over different realizations of the same type of device. The QFlow Lite represents a reference dataset for researchers to use in their experiments for developing QML approaches and concepts. The expanded dataset, denominated QFlow 2.0, consists of $1,599$ idealized simulated measurements stored as $250\times 250$ pixel maps. In addition, the QFlow 2.0 dataset includes two sets of noisy simulated measurements, one with a noise level varied around $1.5$ times the optimized noise level, and the other with a noise level ranging from 00 to $7$ times the optimized noise level.
  • The NTangled Datasets (Schatzki et al., 2021) consist of trained weights for three hardware-efficient Ansatzes), strongly entangling, and convolutional, varying the number of qubits, depths, and types of multipartite entanglement. The states are generated by sending product state inputs into PQCs. The NTangled Datasets are used to benchmark QML models for supervised learning classification tasks.
@@ 6 Tools and Technologies
6.1 Quantum Computing Models
The first concept of a quantum computer was introduced in 1982 by Feynmann (1982). Since then, a significant amount of research efforts have been conducted that resulted in the development of real quantum computers at our disposal for experiments (IBM, b). Due to their different use cases and architectures, quantum computers can be classified into quantum gate models (or quantum circuits) and quantum annealers.
6.1.1 Quantum Gate Model
The gate model involves operations in quantum computation using matrix operators, which are applied to the input qubits. These operators process the information and generate the final output distributions. They are designed with unitary constraints to qualify as valid quantum operators, making them reversible. The reversibility property is necessary for introducing some form of determinism in the quantum computation. A single operator is referred to as a gate in the quantum circuit and can be applied independently or grouped together with other similar quantum operators to perform computation on qubits. A gate can be combined with other gates in varying orientations and series such that they solve certain problems. The collection of gates arranged in configurations respective to the problem being solved along with qubits and measurements is referred to as a circuit (see Section 2.4). Complex problems in the quantum domain can be solved by configuring gates in meaningful ways. The computation done with this quantum model is referred to as the gate model. Since it is easily interpretative and intuitive, it is nowadays the most used quantum computation model.
6.1.2 Quantum Annealer
Quantum Annealing (QA) is an adiabatic quantum computation method that finds optimal solutions to problems. It uses quantum tunneling, entanglement, and superposition to find the solution in terms of the system’s energy. It usually corresponds to finding the global energy minimum in the quantum energy state plane. Instead of using temperature to explore the problem space, QA uses the laws of quantum mechanics to measure the energy state. A quantum annealer is relatively more noise-tolerant compared to the gate model. Quantum annealers shine when they are used to search for the best from many possible combinations. Such challenges are called combinatorial optimization problems (see Section 4.3.4).
@@ 6.2.8 Topological Nanowire Qubits (Majorana Qubits)
Majorana particles are a unique class of topological nanowire particles that are their own antiparticles. The quasi-particles emerge from the topological nature of systems and are bound at zero energy. They are predicted to obey non-abelian braiding statistics such that they can “remember” their histories. A pair of Majorana fermions represents a single qubit, and its realization could lead the way to fault-tolerant quantum computing (Tutschku et al., 2020). Due to their stability at close to normal temperatures, Majorana Qubits work fine and are the technology chosen by Microsoft (Aghaee et al., 2023).

### v4

@@ 5 Datasets
In QML, we distinguish between (i) classical data that must be pre-processed and encoded into quantum states for use in quantum circuits, and (ii) quantum-native data originating from simulations or experiments, which can be ingested more directly by QML models. Below, we separate these concerns into two parts: classical data handling (preprocessing and encoding) and quantum datasets (collections of quantum-native examples and their typical usage).
@@ 5.1 Data Pre-Processing
Since QML has not yet extensively demonstrated its advantages compared to classical ML, it is more susceptible to adoption by the industry. The current quantum computers have few qubits to test and are noisy, making it difficult to demonstrate the current and potential quantum advantage of QML methods. Utilizing quantum methods for ML in the pre-processing phase, it is possible to achieve better classical encoding and performance of quantum classifiers (Mancilla & Pere, 2022).
Since many proposed QML applications rely on using well-known datasets, especially in the context of quantum data with constraints generated due to the underlying quantum mechanics for representing it as quantum states, it is important to formulate and standardize its collection and encoding processes over time to maintain quality (Sierra-Sosa et al., 2021).
@@ 5.2 Data Encoding
5.3 Quantum-Native Data
Some examples of quantum data can be:
  • Discrete data (binary or integer values).
  • Real continuous data (floating-point values).
  • Complex continuous data (complex values).
5.4 Motivation for QML Datasets
By aligning dataset choice with the intended QML task and the available hardware (qubit count, coherence), researchers can fairly benchmark quantum models against classical baselines and progressively build toward demonstrations of quantum advantage.
@@ 5.5 Quantum Datasets Collection
Selecting appropriate datasets for QML is a crucial step that impacts both algorithm design and performance benchmarks. Key considerations include the type of data (quantum vs. classical), the feature dimensionality, and the feasibility of encoding data into quantum states given current hardware constraints. Classical datasets (e.g., images or molecule descriptors) are plentiful and often used, but encoding high-dimensional classical information onto qubits can require significant overhead (one qubit per feature or complex amplitude encoding) and careful embedding schemes. By contrast, quantum-native datasets, i.e., data originating from quantum states or simulations, may better exploit QML’s strengths. Researchers must balance dataset complexity with quantum hardware limits, often resorting to downscaled or synthetic data to fit the available qubit count and circuit depths. Typical benchmark tasks in QML include classification (e.g., distinguishing quantum states or classifying data points), regression (e.g., predicting molecular energies), Hamiltonian learning (inferring system Hamiltonians or optimal controls from data), and state discrimination (distinguishing or identifying quantum states). The choice of dataset should align with the target task: for instance, quantum chemistry datasets enable testing QML for property prediction, whereas quantum state datasets probe a model’s ability to learn inherently quantum patterns. In all cases, benchmarking on well-defined datasets helps compare quantum vs. classical approaches and track quantum advantage. We note that developing standardized quantum datasets (with easily preparable states and known properties) is an active effort to facilitate fair and relevant QML evaluations(Schatzki et al., 2021; Huang et al., 2022).
@@
  • The QDataSet (Perrier et al., 2022), released in 2022, is composed of $52$ datasets based on simulations of one-qubit and two-qubit systems evolving in the presence or absence of noise. It provides a large-scale set of datasets for QML practitioners to train, benchmark, and develop classical and quantum algorithms for common tasks in quantum sciences, such as quantum control, quantum tomography, and noise spectroscopy. It has been generated using customized code running on base-level Python packages to facilitate interoperability and portability across common QML platforms. Each dataset consists of $10,000$ samples and includes a range of information (stored in list, matrix, or tensor format) regarding quantum systems and their evolution, such as quantum state vectors, drift and control Hamiltonians, unitaries, Pauli measurement distributions, time series data, pulse sequence data for square and Gaussian pulses, noise and distortion data. The total compressed size of the QDataSet (compressed with Pickle and zip) is around $14$ TB, while its uncompressed size is around $100$ TB. Researchers use QDataSet to train and evaluate models on tasks like quantum control (designing pulse sequences to reach target states), quantum tomography (reconstructing quantum states from measurement data), and noise spectroscopy (learning noise parameters). A major advantage is its breadth for testing algorithms’ scalability, but its sheer size (terabyte-scale) poses practical loading and memory challenges. Encoding this data into quantum circuits is non-trivial, since each sample may represent a quantum state or time series that would require many qubits or time steps to embed. Thus, QDataSet is often used in classical-quantum hybrid experiments (or simulations) where classical ML models or quantum-inspired algorithms ingest the data, rather than directly preparing each sample on quantum hardware. Nevertheless, it serves as an important benchmark for validating QML methods on tasks intrinsic to quantum systems (e.g., learning system Hamiltonians or optimal controls) in a controlled setting.
  • The QM7, QM8, and QM9 datasets are subsets of the GDB-13 (Blum & Reymond, 2009) and GDB-17 (Ruddigkeit et al., 2012) databases, which contain billions of organic molecules. The QM7 (Rupp et al., 2012) is composed of $7,165$ molecules of up to $23$ atoms (including $7$ heavy atoms such as C, N, O, and S), represented using the Coulomb matrix. An extension of the QM7 dataset for multitask learning, called QM7B (Montavon et al., 2013), has $13$ additional properties (e.g., polarizability, HOMO and LUMO eigenvalues, excitation energies) and contains a total of $7,211$ molecules. The QMB8 (Ramakrishnan et al., 2015) has a training set of $10,000$ molecules with up to $8$ heavy atoms. The QM9 (Ramakrishnan et al., 2014) contains $133,885$ molecules with up to $9$ heavy atoms. In QML, these datasets serve as representative regression benchmarks for molecular properties (e.g., predicting atomization energies. Hybrid quantum-classical models (such as variational quantum algorithms or quantum kernel methods) have been tested on subsets of QM7/QM8/QM9 to evaluate their performance relative to classical models in chemistry tasks. For example, a quantum kernel regression approach was shown to predict QM9 molecular energies with competitive accuracy, hinting that quantum feature maps can capture molecular structure–property relationships (Chang, 2022). However, these datasets pose encoding challenges: each molecule must be described by high-dimensional features. To input them into a quantum model, techniques like dimensionality reduction or efficient feature encoding are required to avoid using hundreds of qubits. The relatively large dataset sizes also mean that current QML experiments typically train on smaller samples or simplified tasks. Even so, QM7/8/9 provide valuable testbeds for QML models aiming to learn complex quantum-mechanical relationships in chemistry, and they are often used to benchmark quantum regression algorithms or to assess quantum models’ ability to generalize from limited training data.
@@
  • The PennyLane Datasets (Bergholm et al., 2018) are composed of the Quantum Chemistry (QChem) Datasets for common molecular systems and the Quantum Many-Body Physics (QMBP) Datasets. The QChem datasets contain the electronic structure data for $42$ different geometries of molecules such as linear hydrogen chains, metallic and non-metallic hydrides, and charged species. For each geometry of the molecule, it is possible to extract molecular data (i.e., information regarding the molecule), Hamiltonian data (i.e., Hamiltonian for the molecular system), tapering data (i.e., features based on Z2 symmetries of the molecular Hamiltonian for performing tapering), tapered observables data, and variational data. The QMBP datasets contain spin models displaying quantum correlations. For each spin system, datasets are generated for $1$-D lattices (linear chain) and $2$-D lattices (rectangular grid). Each dataset contains values for $100$ different configurations of tunable parameters such as the external magnetic field and coupling constants. The PennyLane datasets are commonly used to benchmark variational quantum algorithms and QNNs on physically motivated problems. The QChem datasets provide molecular Hamiltonians and ground state energies, supporting studies where quantum models, such as VQEs, aim to predict molecular energy landscapes across different geometries (Romero et al., 2018; Huggins et al., 2021). The QMBP datasets offer spin system data suitable for phase recognition and classification, allowing quantum classifiers to identify ordered and disordered phases based on Hamiltonian configurations (Cong et al., 2019). Since these datasets are expressed in quantum-native formats (Pauli operators, energy spectra), they can be integrated naturally into quantum circuits. However, encoding continuous variables, such as coupling strengths or bond lengths, remains challenging and often requires data re-uploading techniques (Pérez-Salinas et al., 2020).
  • The TensorFlow Quantum Data builds upon classical datasets such as MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017). It implements the method proposed by Farhi & Neven (2018) to use binary encoding for representing each pixel with a qubit, with the state depending on the pixel’s value. These classification tasks allow direct comparisons between quantum models and classical deep networks on the same data, exploring whether quantum models can achieve similar accuracy with fewer parameters or different scaling. A key challenge here is the scaling of encoding: encoding even moderate-size images quickly becomes intractable. For instance, a full $28\times 28$ image would require $784$ qubits in a straightforward pixel-to-qubit scheme like angle embedding. As a result, researchers either restrict to very low-resolution images or use feature extraction, like Principal Component Analysis (PCA) as in some QML benchmarks (Yu et al., 2019) to reduce dimensionality. While proof-of-concept results like Huang et al. (2021) have shown that quantum models can learn from such data, a concrete quantum advantage may only emerge in the fault-tolerant quantum computing era.
@@
  • The Mendley Datasets for Quantum Circuits Mapping (Acampora et al., 2021) helps to solve the circuit mapping problem as a classification task in a supervised learning setting. Each dataset contains features related to the calibration data of the physical device and others related to the generated quantum circuit for specific IBM quantum machines, such as IBMQ Santiago, IBMQ Athens, and IBMQ 16 Melbourne. By training on these features and labels, classical ML models have been developed to predict good qubit mappings for new circuits, effectively learning to optimize circuit placement without running an exhaustive search. In the context of QML, it is also possible to train a quantum model (e.g., a variational classifier) on this data to evaluate whether a quantum learner can assist in compiler optimization.
@@
  • The QFlow Lite dataset (Zwolak et al., 2018) contains a Python-based software suite to train neural networks to recognize the state of a device and differentiate between states in experimental data. It consists of $1,001$ idealized simulated measurements with gate configurations sampling (stored as $100\times 100$ pixel maps) over different realizations of the same type of device. The QFlow Lite represents a reference dataset for researchers to use in their experiments for developing QML approaches and concepts. The expanded dataset, denominated QFlow 2.0, consists of $1,599$ idealized simulated measurements stored as $250\times 250$ pixel maps. In addition, the QFlow 2.0 dataset includes two sets of noisy simulated measurements, one with a noise level varied around $1.5$ times the optimized noise level, and the other with a noise level ranging from 00 to $7$ times the optimized noise level. For QML research, QFlow Lite offers a chance to test quantum models on tasks of state discrimination in a setting that mimics real quantum sensor data: a quantum model could be trained to recognize charge configurations from the pixelated sensor output. The encoding challenge here is substantial because raw images (often $100\times 100$ or more pixels) would require hundreds of qubits if encoded pixel-by-pixel. Thus, practical QML studies might use downsampled images or extract key features (such as sensor line cuts or PCA components) to feed into a quantum circuit.
  • The NTangled Datasets (Schatzki et al., 2021) consist of trained weights for three hardware-efficient Ansatzes), strongly entangling, and convolutional, varying the number of qubits, depths, and types of multipartite entanglement. The states are generated by sending product state inputs into PQCs. This resource is explicitly used to benchmark quantum classifiers and QNNs on tasks that involve quantum data. A typical use case is a state classification problem: given measurement data (or the parameter vector) from an unknown state, the task is to determine which entanglement class the state belongs to (for instance, distinguish between product states, GHZ-type entangled states, W-type states, etc.). In the NTangled study, the authors trained QML models (like variational quantum classifiers) on these state datasets and evaluated their accuracy in recognizing entanglement patterns (Schatzki et al., 2021). A big advantage of NTangled is that the data is inherently quantum (provided as quantum circuits or state vectors), allowing quantum models to potentially handle it more naturally than classical models. However, using NTangled on real hardware would require preparing those quantum states on a quantum computer and then running the learning algorithm, which is resource-intensive. Overall, NTangled’s primary use is to test and improve QML methods for entanglement recognition and classification, showcasing the promise of quantum-enhanced learning on data that classical methods find challenging.
@@ 6.1.8 Topological Nanowire Qubits (Majorana Qubits)
Majorana particles are a unique class of topological nanowire particles that are their own antiparticles. The quasi-particles emerge from the topological nature of systems and are bound at zero energy. They are predicted to obey non-abelian braiding statistics such that they can “remember” their histories. A pair of Majorana fermions represents a single qubit, and its realization could lead the way to fault-tolerant quantum computing (Tutschku et al., 2020). Despite some initial hurdles (Frolov, 2021; Legg, 2025), Majorana Qubits are extremely promising due to their stability at close to normal temperatures, and are the technology chosen by Microsoft (Aghaee et al., 2023; Aasen et al., 2025).
6.1.9 Pros and Cons of Leading Quantum Technologies
Superconducting Qubits: Pros: Fast gate speeds (nanoseconds) and high gate repetition rates, with solid-state fabrication methods that have already scaled to hundreds of qubits (demonstrated in chips by Google, IBM, etc.), showing compatibility with integrated circuit techniques for scalability. Cons: Relatively short coherence times (typically tens of microseconds up to a few hundred microseconds at best), meaning qubits quickly lose quantum information, and noise from materials and control electronics leads to error rates on the order of $10^{-3}$$10^{-4}$ per gate (Kjaergaard et al., 2020). They also require dilution refrigerators operating at millikelvin temperatures for stability, incurring significant overhead in operational temperature and complexity.
Trapped Ion Qubits: Pros: Individual ions confined in electromagnetic traps have very long coherence times (on the order of seconds or more, as the ions’ internal states are well-isolated from the environment). They achieve high-fidelity gates (single- and two-qubit gate errors as low as $10^{-3}$ or better) and all-to-all connectivity within a trap, making noise resilience and entangling operations robust (Bruzewicz et al., 2019). Trapped ions operate in ultra-high vacuum but essentially at room temperature (no cryogenics needed, only laser cooling), greatly easing the operational requirements. Cons: Gate speeds are much slower (two-qubit gate durations in the millisecond range), limiting the processing speed. Scaling to large numbers of qubits in one device is challenging. Ion traps with more than 50 qubits face mode crowding and crosstalk, and while modular architectures (networking multiple traps) are proposed, they add complexity. Thus, scalability is moderate, since trapped-ion systems excel in qubit quality but require significant engineering (laser systems, stable trapping) to handle many qubits and fast operations.
Neutral Atom Qubits: Pros: Neutral atoms (e.g., atoms in optical tweezers) offer inherently scalable architectures. Hundreds of atoms can be arranged in 1D or 2D arrays with each atom as an identical qubit. They have long coherence times using stable hyperfine states (similar to trapped ions), and use highly excited Rydberg states to induce strong, controllable interactions for two-qubit gates (enabling entanglement across neighboring atoms). Neutral atom platforms operate at near-room temperature (cold atoms are trapped by lasers in vacuum, but no dilution fridge is required). Cons: Two-qubit gate fidelities, while improving, are currently lower than in ions or superconductors (errors on the order of a few percent), due to relative decoherence of short-lived Rydberg excitations and sensitivity to laser phase noise (Henriet et al., 2020; Shi, 2022). Achieving uniform control laser focus and stability across a large array is technically complex, which can impact scalability even though the physical layout is extensible. Additionally, gate speeds for Rydberg-mediated operations are intermediate (microsecond scale), so while faster than ions, some algorithms still face time limitations given qubit coherence constraints. Overall, neutral atoms promise high qubit counts and reconfigurability, but require further advances in control precision and error correction to match the gate fidelities of other technologies.
Photonic Qubits: Pros: Photonic quantum computing uses photons (light) as qubits, which naturally suffer virtually no decoherence, because a photon traveling through space maintains its quantum state until interaction or loss, giving effectively infinite coherence in transit. Photonic systems can operate at ambient conditions (operational temperature is typically room temperature), and are ideal for communication (flying qubits) and networking of quantum information. They also offer large Hilbert spaces (e.g., encoding in light modes allows many-state qubit generalizations). Cons: Direct two-qubit interactions between photons are difficult; operations often rely on probabilistic gates or effective interactions mediated by matter, which significantly complicates scalability for quantum computing purposes. Scaling photonic qubit counts and circuits is challenging due to losses. As the number of components (beam splitters, phase shifters, detectors) grows, photon loss and detection inefficiency cause error rates to rise. Noise resilience in terms of decoherence is high, but photon loss acts as a dominant noise source. Moreover, while photons do not need cooling to preserve state, high-efficiency single-photon sources and detectors often require cryogenic technology, adding to the experimental overhead (Slussarenko & Pryde, 2019). In summary, photonic qubits excel in coherence and connectivity (they can be sent over long distances), but the difficulty of implementing deterministic entangling gates and the challenges in reliably generating and detecting large numbers of photons make building large-scale photonic quantum computers an ongoing research effort.
@@ 6.3 Quantum Simulators and Primitives
6.3.1 Comparison of Quantum Software Frameworks
While all the above frameworks enable quantum programming, each has distinct strengths and target use cases. IBM’s Qiskit is a comprehensive SDK spanning low-level pulse control to high-level algorithms, with extensive libraries (for chemistry, finance, machine learning, etc.) and a large community. It is well-suited for general-purpose quantum computing across a broad user base, from beginners to advanced researchers (Qiskit contributors, 2023). In contrast, Google’s Cirq focuses on giving experts fine-grained control over quantum circuits and is tailored to NISQ devices where hardware details matter for performance (e.g., mapping to qubit topology, scheduling gates with minimal error) (Developers, 2023). Cirq is often chosen for research that directly interfaces with hardware (such as calibrations or implementing algorithms on Google’s Sycamore processor), whereas Qiskit emphasizes a more hardware-agnostic approach (with cloud access to IBM quantum processors) and a rich set of built-in algorithmic modules. Rigetti’s PyQuil lies in between: it is designed for writing programs in the Quil assembly language to run on Rigetti’s superconducting qubit hardware (or its Quantum Virtual Machine simulator), allowing explicit hybrid quantum-classical computations with shared memory (Smith et al., 2016). This makes PyQuil advantageous for experiments requiring classical feedback within quantum circuits (a feature now emerging in other frameworks as well), though its ecosystem is more tied to Rigetti-specific tools and hardware. Finally, Xanadu’s PennyLane sets itself apart by targeting QML and differentiable programming. It provides automatic differentiation of quantum circuits and integration with popular ML libraries, enabling optimization of quantum circuits in conjunction with classical neural networks (Bergholm et al., 2018). PennyLane is hardware-agnostic via plugin interfaces, as users can prototype algorithms on simulators or real devices from IBM, Google, Honeywell, etc., through a unified high-level interface.
In summary, Qiskit and Cirq are often favored for developing quantum algorithms and exploring hardware-specific optimizations (with Qiskit offering a broader platform compatibility and Cirq a lower-level control), PyQuil is specialized for co-processing with Rigetti QPUs using a unique ISA, and PennyLane excels in hybrid quantum-classical workflows, particularly for variational algorithms and quantum neural networks. Each framework’s design reflects its priorities: Qiskit for completeness and accessibility, Cirq for hardware-centric research, PyQuil for integrating quantum programs with classical computation, and PennyLane for machine-learning-driven quantum experimentation.